OpenStudy (anonymous):

Alex and 3 friends are going Roller Skating. It costs $7.25 for admission and she is buying each friend a Coke for $1.75. How much will it cost for Alex to pay for everyone in her group? Write a number sentence to show how you computed it Alex and 3 friends are going Roller Skating. It costs $7.25 for admission and she is buying each friend a Coke for $1.75. How much will it cost for Alex to pay for everyone in her group? Write a number sentence to show how you computed it @Mathematics

OpenStudy (anonymous):

is she included in the everyone?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you are supposed to multiply 4 (including her) time 7.25. And then you can multiply 1.75 times 4 (including her) and add up the products of both the answers. so it would be 7.25*4=29 and 1.75*4=7. You add the 29+7 and get 36.

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

wait but the thing im confused on is what would be the number sentance ?

OpenStudy (anonymous):

it would be 7.25*4=29 and the second would be 1.75*4=7, and the last would be 29+7=36. I suggest you use all three because it shows how you got your answer.

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

kinda but how would i put those 3 together?

OpenStudy (anonymous):

ok so what you do is 7.25*4+1.75*4=36..do you get why though?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

ok good luck!

OpenStudy (anonymous):

thanks